OpenStudy (anonymous):

find exponential equation e^2x-4^ex+3=0

OpenStudy (anonymous):

is that the equation? it doesn't look right

myininaya (myininaya):

Do you know how to solve \[u^2-4u+3=0\]

myininaya (myininaya):

hey ninja i assume the person meant \[e^{2x}-4e^{x}+4=0\] \[(e^x)^2-4e^x+4=0\] All I did was replace e^x with u

OpenStudy (anonymous):

e^2x-4^ex+3=0 Let x = e^x: X^2 -4X + 3 = 0 -> X = 1, X = 3 Thus e^x =1, 3 Take Ln: x = ln1 = 0 X = ln3 = 1.1